a plane is flying horizontally at 98m/s and releases d object which reaches the ground in 10 sec the angle made by packet while hitting ground is

Grade:12th pass

2 Answers

Indrajit dodiya
6 years ago
In 10 second horizontal distance travelled by plane is 980 m and vertical distance is 490 m so angle is tan^-1 (0.5)
DIWAKAR PANDEY
6 years ago
A plane is moving with 98m/s velocity.An object is released from it so it have acceleration of "g". So g=v/t so velocity of object will be V=gt V=9.8×10=98m/s Now using tan x = 98/98=1 So value of tan x =1 So value of tanx is 1when X=45°